OpenStudy (anonymous):

Solve and write your solution in interval notation..\[-3\le2x-1<5\]

OpenStudy (anonymous):

-2 <= 2x < 6 -1 <= x < 3

OpenStudy (anonymous):

1) add \(1\) all the way across 2) divide everything by \(2\)

OpenStudy (anonymous):

^ what the bicycle said

OpenStudy (anonymous):

I got \[-3\le2x-1<5=-3+1\le2x+1<5+1=-4\le2x <6 =-2lex <3\]

OpenStudy (anonymous):

A little mistake, -3 + 1 = -2...you wrote -4

OpenStudy (anonymous):

Oops..yes you are right@ Cormacpayne

OpenStudy (anonymous):

how do I write it in interval notation? \[(-1,3\right]\]

OpenStudy (anonymous):

lol \[[-1,3)\] i think is what you meant

OpenStudy (anonymous):

^ that

OpenStudy (anonymous):

which is the same as -1 <= x < 3